# DeepResonance: Enhancing Multimodal Music Understanding via Music-centric Multi-way Instruction Tuning

### v1

@@ Abstract
Recent advancements in music large language models (LLMs) have significantly improved music understanding tasks, which involve the model’s ability to analyze and interpret various musical elements. These improvements primarily focused on integrating both music and text inputs. However, the potential of incorporating additional modalities such as images, videos and textual music features to enhance music understanding remains unexplored. To bridge this gap, we propose DeepResonance, a multimodal music understanding LLM fine-tuned via multi-way instruction tuning with multi-way aligned music, text, image, and video data. To this end, we construct Music4way-MI2T, Music4way-MV2T, and Music4way-Any2T, three 4-way training and evaluation datasets designed to enable DeepResonance to integrate both visual and textual music feature content. We also introduce multi-sampled ImageBind embeddings and a pre-alignment Transformer to enhance modality fusion prior to input into text LLMs, tailoring DeepResonance for multi-way instruction tuning. Our model achieves state-of-the-art performances across six music understanding tasks, highlighting the benefits of the auxiliary modalities and the structural superiority of DeepResonance. We plan to open-source the models and the newly constructed datasets.
@@ 1 Introduction
We present DeepResonance, a multimodal music understanding LLM trained on music-centric multi-way data that integrates the music, text, image, and video modalities. Specifically, we construct two 4-way training datasets, Music4way-MI2T and Music4way-MV2T, where the source data comprises music, images/videos, and textual descriptions and instructions. The target data includes textual descriptions enriched with multimodal information, such as music, images/videos, and low-level music features, including tempo, chords, key, and downbeats. Using these datasets, we develop DeepResonance based on the NExT-GPT Wu et al. (2024) architecture. To enhance multimodal music understanding tasks, we propose two key modifications to the backbone model. The first involves multi-sampled ImageBind embeddings, which are designed to retain richer information of music, image, and video modalities from the ImageBind encoders Girdhar et al. (2023), thereby fostering deeper interaction with the music modality and improving multimodal music understanding. The second is a pre-alignment Transformer, a module aimed at pre-adapting different modalities to each other before feeding them into the text LLM module. This component is particularly effective in simultaneously handling multimodal inputs of music, text, images, and videos. These innovations collectively advance the model’s ability to integrate and process diverse multimodal signals for enhanced music understanding.
@@
  • We propose multi-sampled ImageBind embeddings and a pre-alignment Transformer to enhance multimodal fusion for music LLMs.
@@
Figure 2: Overview of DeepResonance. It integrates the “multi-way unified caption” as a target, along with “multi-sampled ImageBind embeddings” and a “pre-alignment Transformer” as novel architectural components.
@@ 2 Related Work
Our work differs from the previous studies on music understanding and music foundation models in three aspects: (1) We are the first to apply instruction-tuning on music foundation models using multi-way data that integrates multiple modalities, shifting the paradigm from the traditional music + text $\xrightarrow{}$ text approach to the music + image/video + text $\xrightarrow{}$ text (multimodal-enriched) framework. (2) We are the first to access the generalization of music LLMs to multi-way multimodal inputs using our newly curated Music4way-Any2T dataset and zero-shot evaluation on out-of-domain benchmarks. (3) We propose multi-sampled ImageBind embeddings and pre-alignment Transformer that significantly impact multimodal music understanding tasks, delivering state-of-the-art music LLMs optimized for different downstream tasks.
@@ 3.1 Music-centric Multi-way Datasets Construction
Music4way: Building upon the AudioSet Gemmeke et al. (2017) music clips filtered by M2UGen [^1] https://github.com/shansongliu/M2UGen, we curate a multi-way aligned dataset, comprising a total of 172.57 hours of music. [^2] Note that in M2UGen, each music clip is paired with only one other modality during training. Each video-music pair from AudioSet is processed as shown in Fig. 3 (left). First, we randomly extract a video frame to serve as the aligned image for the video and music. Next, BLIP Li et al. (2022), VideoMAE Tong et al. (2022), and MU-LLaMA Liu et al. (2024) are used to generate captions for the image, video, and music, respectively. Following LLark Gardner et al. (2024), we employ madmom Böck et al. (2016) to extract low-level music features, including tempo, chords, downbeats, and key, representing these features in textual form. This process creates a 4-way alignment for each music-video pair, establishing relationships between music, image, video, and text. The text modality conveys information through captions or feature values derived from the other three modalities. The resulting Music4way dataset includes 59,128 training samples (164.24 hours) and 3,000 evaluation samples (8.33 hours).
@@ 3.2 Multi-way Instruction Tuning Data Construction
Music4way-MI2T and Music4way-MV2T: Inspired by VAST Chen et al. (2023), which established multimodal connections between vision, audio, and subtitles to generate omni-modal captions for improving visual-text tasks, we extend this approach to the music domain, exploring its potential benefits for music understanding tasks. As shown in Fig. 3, we prompt the open-source MPT-7B Team (2023) model to combine captions from images, videos, and music, as well as textual music features, creating a multi-way unified caption. These captions are then paired with music and image/video inputs to form the source-target pairs required for fine-tuning LLMs. For the task of music + image to multi-way unified caption, we construct the Music4way-MI2T dataset. Similarly, for music + video inputs, we construct the Music4way-MV2T dataset. These datasets are designed to train the music understanding model to infer visual content and musical details (e.g., instruments, sentiments, and low-level features) directly from raw music and image/video files. The templates we utilized for prompting MPT-7B to generate the multi-way unified captions are provided in the Appx. C. As shown in Fig. 2, these datasets are then used to fine-tune music understanding models, with inputs comprising music and image/video, a fixed instruction template (see Appx. C), and outputs of multi-way unified captions. Additionally, we create corresponding test sets from the test split of Music4way to benchmark performance on multimodal music understanding tasks. (See Sec. 5.3)
Music4way-Any2T: To evaluate the robustness of multimodal music LLMs to diverse text inputs and queries, we introduce the Music4way-Any2T dataset. As shown in Fig. 3, this dataset features flexible inputs, allowing each modality to appear in any position. It includes instructions and outputs presented as diverse question-answer pairs, covering various aspects (e.g., visual or musical content) of the multi-way aligned data. To construct structured data with input, instruction, and output fields, we use GPT-4o mini [^3] https://platform.openai.com/docs/models, prompted with all textual information from the Music4way dataset alongside the multi-way unified caption. Detailed prompting templates and data examples can be found in the Appx. D. The Music4way-Any2T dataset is used to benchmark the robustness and generalization capabilities of multimodal music understanding models, with results presented in Sec. 5.3.
@@ 4 Model Architecture Tailored for Multimodal Music Understanding
To address these challenges, we propose two modules for music LLMs: multi-sampled ImageBind embeddings and pre-alignment Transformer, as shown in Fig. 2. The former leverages embeddings from multiple clips sampled by ImageBind without pooling, while the latter introduces a Transformer to globally integrate and align information across modalities before feeding it into the LLM. Formally, the proposed model is defined as
@@
Here, $\mathcal{T}$ represents the pre-alignment Transformer, and $\mathbf{e}_{*}$ ($*\in\{m,v,i\}$) from Eq. 1 is reformulated as $\{\mathbf{e}_{*}\}_{1:\mathrm{N}_{*}}$ to incorporate multi-sampled ImageBind embeddings for each modality, where $\mathrm{N}_{*}$ denotes the number of sampled clips for a given modality. With these components, the multi-sampled ImageBind embeddings preserve finer-grained information for each modality, which is expected to enhance music understanding tasks (Sec. 5). The pre-alignment Transformer uses bidirectional attention to encode cross-modal dependencies, enhancing the interactions between music and other modalities, ultimately aiming to improve multimodal music understanding tasks (Sec. 5.3).
@@ B-1 R-L BERTS B-1 R-L BERTS B-1 R-L BERTS
Table 2: Results on MusicQA, MusicCaps, and Music4way-MusicCaps. The top two performances are highlighted in bold. “*” denotes the test data was included in the corresponding model’s training set. “” indicates that the test set’s train split is included in the corresponding model’s training data, representing a supervised setting. “B-1”, “R-L”, and “BERTS” denote BLEU-1, ROUGE-L, and BERTScore, respectively.
@@ 5 Experiments and Results
Following the training strategy of NExT-GPT and M2UGen, we train DeepResonance in two stages. Table 1 summarizes all datasets used for training, with training details provided in Appx. B. Subsequently, we evaluate DeepResonance across three music understanding tasks and three multimodal music understanding tasks in supervised settings. Additionally, we assess the model’s zero-shot performance on out-of-domain datasets and conduct an ablation study to demonstrate the effectiveness of each proposed component. Results are reported using BLEU Papineni et al. (2002), [^4] Following prior studies, we primarily report BLEU-1, while BLEU-1 and BLEU are shown in Appx. EFG, and H. ROUGE-L Lin (2004), and BERTScore Zhang et al. (2020).
@@ 5.1 Baselines and Ours
Below are the baseline models and our models that we compare in this section.
@@
DeepResonance (Ours): The models introduced in Sec. 4, built upon NExT-GPT and trained with datasets as shown in Table 1. We introduce two DeepResonance variants: DeepResonance-$\alpha$, trained without the pre-alignment Transformer, and DeepResonance-$\beta$, which includes it. The following sections examine their effectiveness, as the pre-alignment Transformer is specifically designed for multimodal music understanding tasks (Sec. 5.3).
@@ B-1 R-L BERTS B-1 R-L BERTS B-1 R-L BERTS
DeepResonance-$\alpha$ (ours) 44.6 30.1 89.4 44.5 29.6 89.2 54.7 35.1 90.9
@@ 5.2 Music Understanding Tasks
First, we observe that DeepResonance outperforms baseline models on three datasets, demonstrating the effectiveness of our proposed training datasets and model architecture. Second, we find that applying the same training data to the NExT-GPT framework yields a better performance than using the M2UGen framework, suggesting that NExT-GPT is a more suitable backbone model. Third, by expanding M2UGen’s training data with our constructed Music4way dataset, we observe an improvement in performance, further validating the effectiveness of the Music4way dataset. Finally, the comparison between DeepResonance-$\alpha$ and DeepResonance-$\beta$ shows a comparable performance, indicating that the pre-alignment Transformer is not crucial for music + text $\xrightarrow{}$ text tasks. This aligns with expectations, as the pre-alignment Transformer is designed for multimodal music understanding, whereas these tasks involve only a single modality (music) alongside the text query.
Figure 5: Ablation study on Music4way-MI2T, Music4way-MV2T, and Music4way-Any2T. “PT-*L” indicates the number of layers used in the pre-alignment Transformer. See Appx. H for result details.
Figure 6: Ablation study on MusicQA, MusicCaps, and Music4way-MusicCaps. See Appx. H for result details.
@@ 5.3 Multimodal Music Understanding Tasks
First, we observe that Music4way-MI2T and Music4way-MV2T represent novel downstream tasks for which no existing baseline models are inherently equipped. Through supervised fine-tuning with our curated training data for these datasets, the DeepResonance models gain the ability to generate unified multimodal captions successfully. Second, on Music4way-Any2T, which features flexible inputs and open-ended question-answer pairs, the baseline models perform poorly. Their generalization remains weaker than DeepResonance models, highlighting their limitations in handling diverse input patterns. Third, comparing DeepResonance-$\alpha$ and DeepResonance-$\beta$, we find that the latter demonstrates a superior performance on Music4way-MI2T and Music4way-MV2T. This indicates that the pre-alignment Transformer, with its additional parameters, effectively integrates multiple modalities, thereby improving the supervised performance in structured multimodal music understanding tasks. However, DeepResonance-$\beta$ exhibits reduced robustness on Music4way-Any2T, which highlights a trade-off, as the model’s increased complexity may hinder its adaptability with limited instruction tuning data.
@@ 5.4 Zero-shot Evaluation
Fig. 4 present the zero-shot performance of DeepResonance and baselines on music understanding benchmarks, including GTZAN Tzanetakis (2001), MusicNet Thickstun et al. (2017), MTG-Jamendo Bogdanov et al. (2019), MusicInstruct-short, and MusicInstruct-long Deng et al. (2024). We adopt the benchmark settings outlined in OpenMU-bench Zhao et al. (2024), combining captioning and reasoning test sets where separate splits exist. All NExT-GPT-based models without exposure to test data during training are compared, to ensure the zero-shot configurations. Detailed results are provided in the Appx. G.
First, DeepResonance-$\beta$ achieves the best performance across all five benchmarks in terms of ROUGE-L and BERTScore, demonstrating the effectiveness of the proposed training data and model architecture. Second, DeepResonance-$\beta$ outperforms DeepResonance-$\alpha$, highlighting the pre-alignment Transformer’s effectiveness in improving inference on unseen data. Referring back to Sec. 5.3, we recommend DeepResonance-$\alpha$ for tasks with flexible inputs and DeepResonance-$\beta$ for other scenarios. Finally, DeepResonance’s zero-shot performance approaches the supervised upper bounds of MusiLingo on MusicInstruct in terms of BERTScore, demonstrating the out-of-domain generalization capabilities of DeepResonance.

### v2

@@ Abstract
Recent advancements in music large language models (LLMs) have significantly improved music understanding tasks, which involve the model’s ability to analyze and interpret various musical elements. These improvements primarily focused on integrating both music and text inputs. However, the potential of incorporating additional modalities such as images, videos and textual music features to enhance music understanding remains unexplored. To bridge this gap, we propose DeepResonance, a multimodal music understanding LLM fine-tuned via multi-way instruction tuning with multi-way aligned music, text, image, and video data. To this end, we construct Music4way-MI2T, Music4way-MV2T, and Music4way-Any2T, three 4-way training and evaluation datasets designed to enable DeepResonance to integrate both visual and textual music feature content. We also introduce multi-sampled ImageBind embeddings and a pre-LLM fusion Transformer to enhance modality fusion prior to input into text LLMs, tailoring DeepResonance for multi-way instruction tuning. Our model achieves state-of-the-art performances across six music understanding tasks, highlighting the benefits of the auxiliary modalities and the structural superiority of DeepResonance. We plan to open-source the models and the newly constructed datasets.
@@ 1 Introduction
We present DeepResonance, a multimodal music understanding LLM trained on music-centric multi-way data that integrates the music, text, image, and video modalities. Specifically, we construct two 4-way training datasets, Music4way-MI2T and Music4way-MV2T, where the source data comprises music, images/videos, and textual descriptions and instructions. The target data includes textual descriptions enriched with multimodal information, such as music, images/videos, and low-level music features, including tempo, chords, key, and downbeats. Using these datasets, we develop DeepResonance based on the NExT-GPT Wu et al. (2024) architecture. To enhance multimodal music understanding tasks, we propose two key modifications to the backbone model. The first involves multi-sampled ImageBind embeddings, which are designed to retain richer information of music, image, and video modalities from the ImageBind encoders Girdhar et al. (2023), thereby fostering deeper interaction with the music modality and improving multimodal music understanding. The second is a pre-LLM fusion Transformer, a module aimed at pre-adapting different modalities to each other before feeding them into the text LLM module. This component is particularly effective in simultaneously handling multimodal inputs of music, text, images, and videos. These innovations collectively advance the model’s ability to integrate and process diverse multimodal signals for enhanced music understanding.
@@
  • We propose multi-sampled ImageBind embeddings and a pre-LLM fusion Transformer to enhance multimodal fusion for music LLMs.
@@
Figure 2: Overview of DeepResonance. It integrates the “multi-way unified caption” as a target, along with “multi-sampled ImageBind embeddings” and a “pre-LLM fusion Transformer” as novel architectural components.
@@ 2 Related Work
Our work differs from the previous studies on music understanding and music foundation models in three aspects: (1) We are the first to apply instruction-tuning on music foundation models using multi-way data that integrates multiple modalities, shifting the paradigm from the traditional music + text $\xrightarrow{}$ text approach to the music + image/video + text $\xrightarrow{}$ text (multimodal-enriched) framework. (2) We are the first to access the generalization of music LLMs to multi-way multimodal inputs using our newly curated Music4way-Any2T dataset and zero-shot evaluation on out-of-domain benchmarks. (3) We propose multi-sampled ImageBind embeddings and pre-LLM fusion Transformer that significantly impact multimodal music understanding tasks, delivering state-of-the-art music LLMs optimized for different downstream tasks.
@@ 3.1 Music-centric Multi-way Datasets Construction
Music4way: Building upon the AudioSet Gemmeke et al. (2017) music clips filtered by M2UGen, we curate a multi-way aligned dataset, comprising a total of 172.57 hours of music. [^1] Note that in M2UGen, each music clip is paired with only one other modality during training. Each video-music pair from AudioSet is processed as shown in Fig. 3 (left). First, we randomly extract a video frame to serve as the aligned image for the video and music. Next, BLIP Li et al. (2022), VideoMAE Tong et al. (2022), and MU-LLaMA Liu et al. (2024) are used to generate captions for the image, video, and music, respectively. Following LLark Gardner et al. (2024), we employ madmom Böck et al. (2016) to extract low-level music features, including tempo, chords, downbeats, and key, representing these features in textual form. This process creates a 4-way alignment for each music-video pair, establishing relationships between music, image, video, and text. The text modality conveys information through captions or feature values derived from the other three modalities. The resulting Music4way dataset includes 59,128 training samples (164.24 hours) and 3,000 evaluation samples (8.33 hours).
@@ 3.2 Multi-way Instruction Tuning Data Construction
Music4way-MI2T and Music4way-MV2T: Inspired by VAST Chen et al. (2023), which established multimodal connections between vision, audio, and subtitles to generate omni-modal captions for improving visual-text tasks, we extend this approach to the music domain, exploring its potential benefits for music understanding tasks. As shown in Fig. 3, we prompt the open-source MPT-7B Team (2023) model to combine captions from images, videos, and music, as well as textual music features, creating a multi-way unified caption. These captions are then paired with music and image/video inputs to form the source-target pairs required for fine-tuning LLMs. For the task of music + image to multi-way unified caption, we construct the Music4way-MI2T dataset. Similarly, for music + video inputs, we construct the Music4way-MV2T dataset. These datasets are designed to train the music understanding model to infer visual content and musical details (e.g., instruments, sentiments, and low-level features) directly from raw music and image/video files. The templates we utilized for prompting MPT-7B to generate the multi-way unified captions are provided in the Appx. E. As shown in Fig. 2, these datasets are then used to fine-tune music understanding models, with inputs comprising music and image/video, a fixed instruction template (see Appx. E), and outputs of multi-way unified captions. Additionally, we create corresponding test sets from the test split of Music4way to benchmark performance on multimodal music understanding tasks. (See Sec. 5.3)
Music4way-Any2T: To evaluate the robustness of multimodal music LLMs to diverse text inputs and queries, we introduce the Music4way-Any2T dataset. As shown in Fig. 3, this dataset features flexible inputs, allowing each modality to appear in any position. It includes instructions and outputs presented as diverse question-answer pairs, covering various aspects (e.g., visual or musical content) of the multi-way aligned data. To construct structured data with input, instruction, and output fields, we use GPT-4o mini [^2] https://platform.openai.com/docs/models, prompted with all textual information from the Music4way dataset alongside the multi-way unified caption. Detailed prompting templates and data examples can be found in the Appx. F. The Music4way-Any2T dataset is used to benchmark the robustness and generalization capabilities of multimodal music understanding models, with results presented in Sec. 5.3.
@@ 4 Model Architecture Tailored for Multimodal Music Understanding
To address these challenges, we propose two modules for music LLMs: multi-sampled ImageBind embeddings and pre-LLM fusion Transformer, as shown in Fig. 2. The former leverages embeddings from multiple clips sampled by ImageBind without pooling, while the latter introduces a Transformer to globally integrate and align information across modalities before feeding it into the LLM. Formally, the proposed model is defined as
@@
Here, $\mathcal{T}$ represents the pre-LLM fusion Transformer, and $\mathbf{e}_{*}$ ($*\in\{m,v,i\}$) from Eq. 1 is reformulated as $\{\mathbf{e}_{*}\}_{1:\mathrm{N}_{*}}$ to incorporate multi-sampled ImageBind embeddings for each modality, where $\mathrm{N}_{*}$ denotes the number of sampled clips for a given modality. With these components, the multi-sampled ImageBind embeddings preserve finer-grained information for each modality, which is expected to enhance music understanding tasks (Sec. 5). The pre-LLM fusion Transformer uses bidirectional attention to encode cross-modal dependencies, enhancing the interactions across music and other modalities, aiming to improve multimodal music understanding tasks (Sec. 5.3).
@@ B-1 R-L BERTS B-1 R-L BERTS B-1 R-L BERTS
Table 2: Results on MusicQA, MusicCaps, and Music4way-MusicCaps. The top two performances are highlighted in bold. “*” denotes the test data was included in the corresponding model’s training set. “” indicates supervised settings. “B-1”, “R-L”, and “BERTS” denote BLEU-1, ROUGE-L, and BERTScore, respectively.
@@ 5 Experiments and Results
Following the training strategy of NExT-GPT and M2UGen, we train DeepResonance in two stages. Table 1 summarizes all datasets used for training, with training details provided in Appx. B. Subsequently, we evaluate DeepResonance across three music understanding tasks and three multimodal music understanding tasks in supervised settings. Additionally, we assess the model’s zero-shot performance on out-of-domain datasets and conduct an ablation study to demonstrate the effectiveness of each proposed component. Results are reported using BLEU Papineni et al. (2002), [^3] Following prior studies, we primarily report BLEU-1, while BLEU-1 and BLEU are shown in Appx. GHI, and J. ROUGE-L Lin (2004), and BERTScore Zhang et al. (2020).
@@ 5.1 Baselines and Ours
Below are the baselines and ours that we compare.
Existing baseline models: SALMONN Tang et al. (2024a), MU-LLaMA Liu et al. (2024), NExT-GPT Wu et al. (2024), M2UGen Hussain et al. (2023), MusiLingo Deng et al. (2024) and OpenMU Zhao et al. (2024) (details in Appx. C).
@@
DeepResonance (Ours): The models introduced in Sec. 4, built upon NExT-GPT and trained with datasets as shown in Table 1. We introduce two DeepResonance variants: DeepResonance-$\alpha$, trained without the pre-LLM fusion Transformer, and DeepResonance-$\beta$, which includes it. The following sections examine their effectiveness, as the pre-LLM fusion Transformer is specifically designed for multimodal music understanding tasks (Sec. 5.3). We also present the performance of DeepResonance models trained using the Mistral- and LLaMA-3-based NExT-GPT in Appx. G.
@@ B-1 R-L BERTS B-1 R-L BERTS B-1 R-L BERTS
DeepResonance-$\alpha$ (ours) 48.7 36.2 90.2 48.9 36.5 90.2 37.2 29.6 89.5
@@ 5.2 Music Understanding Tasks
First, we observe that DeepResonance outperforms baseline models on three datasets, demonstrating the effectiveness of our proposed training datasets and model architecture. Second, we find that applying the same training data to the NExT-GPT framework yields a better performance than using the M2UGen framework, suggesting that NExT-GPT is a more suitable backbone model. Third, by expanding M2UGen’s training data with our constructed Music4way dataset, we observe an improvement in performance, further validating the effectiveness of the Music4way dataset. Finally, the comparison between DeepResonance-$\alpha$ and DeepResonance-$\beta$ shows a comparable performance, indicating that the pre-LLM fusion Transformer is not crucial for music + text $\xrightarrow{}$ text tasks. This aligns with expectations, as the pre-LLM fusion Transformer is designed for multimodal music understanding, whereas these tasks involve only a single modality (music) alongside the text query.
Figure 5: Ablation study on Music4way-MI2T, Music4way-MV2T, and Music4way-Any2T. “PT-*L” indicates the number of layers used in the pre-LLM fusion Transformer. See Appx. J for result details.
Figure 6: Ablation study on MusicQA, MusicCaps, and Music4way-MusicCaps. See Appx. J for result details.
@@ 5.3 Multimodal Music Understanding Tasks
First, we observe that Music4way-MI2T and Music4way-MV2T represent novel downstream tasks for which no existing baseline models are inherently equipped. Through supervised fine-tuning with our curated training data for these datasets, the DeepResonance models gain the ability to generate unified multimodal captions successfully. Second, on Music4way-Any2T, which features flexible inputs and open-ended question-answer pairs, the baseline models perform poorly. Their generalization remains weaker than DeepResonance models, highlighting their limitations in handling diverse input patterns. Third, comparing DeepResonance-$\alpha$ and DeepResonance-$\beta$, we find that the latter demonstrates a superior performance on Music4way-MI2T and Music4way-MV2T. This indicates that the pre-LLM fusion Transformer, with its additional parameters, effectively integrates multiple modalities, thereby improving the supervised performance in structured multimodal music understanding tasks. However, DeepResonance-$\beta$ exhibits reduced robustness on Music4way-Any2T, which highlights a trade-off, as the model’s increased complexity may hinder its adaptability with limited instruction tuning data.
@@ 5.4 Zero-shot Evaluation
Fig. 4 present the zero-shot performance of DeepResonance and baselines on music understanding benchmarks, including GTZAN Tzanetakis (2001), MusicNet Thickstun et al. (2017), MTG-Jamendo Bogdanov et al. (2019), MusicInstruct-short, and MusicInstruct-long Deng et al. (2024). We adopt the benchmark settings outlined in OpenMU-bench Zhao et al. (2024), combining captioning and reasoning test sets where separate splits exist. All NExT-GPT-based models without exposure to test data during training are compared, to ensure the zero-shot configurations. Detailed results are provided in the Appx. I.
First, DeepResonance-$\beta$ achieves the best performance across all five benchmarks in terms of ROUGE-L and BERTScore, demonstrating the effectiveness of the proposed training data and model architecture. Second, DeepResonance-$\beta$ outperforms DeepResonance-$\alpha$, highlighting the pre-LLM fusion Transformer’s effectiveness in improving inference on unseen data. Referring back to Sec. 5.3, we recommend DeepResonance-$\alpha$ for tasks with flexible inputs and DeepResonance-$\beta$ for other scenarios. Finally, DeepResonance’s zero-shot performance approaches the supervised upper bounds of MusiLingo on MusicInstruct in terms of BERTScore, demonstrating the out-of-domain generalization capabilities of DeepResonance.